OpenStudy (anonymous):

How do i show this? \[\left( a X b \right) Dot b = 0\] for all vectors a and b in V3

myininaya (myininaya):

maybe we can do something with this i haven't decided yet

myininaya (myininaya):

and then we can write x(bz-yc+y^2-zb)+y(-az+az)=0 if i didn't make a mistake x(-yc+y^2)+y(0)=0 x(-yc+y^2)=0 xy(y-c)=0

myininaya (myininaya):

x=0, y=0, y=c (this should imply c=0 since y=0)

myininaya (myininaya):

so we have vector A=<a,b,0> and B=<0,0,z> what do you think?

myininaya (myininaya):

anything in this form as long as you have x=y=c=0

myininaya (myininaya):

i think i like this problem